OpenStudy (anonymous):

Write an equation of the line, in point-slope form, that passes through the two given points. points: (–17, 8), (3, —2) (y + 8) = –2(x –17) (y – 8) = -1/2(x + 17) (y – 8) = –2(x +17) (y + 8) = -1/2(x – 17)

OpenStudy (anonymous):

y-8=-2(x+17)

OpenStudy (anonymous):

well you are given two points so you can find your slope which is m= y2-y1 / x2-x1 so you get -2-8/ 3-(-17) = -1/2

OpenStudy (anonymous):

First find slope by two points y2-y1/x2-x1.Then Use Point Slope Form y-y1=m(x-x1)

OpenStudy (anonymous):

So , it would be the 2nd one ?

OpenStudy (anonymous):

nooo , the 4th one i mean . . .

OpenStudy (anonymous):

point-slope form is y-y1=m(x-x1) so you can use one point any of your choice . im going to use the point (3,-2), and you get y-(-2) = -1/2 (x-3)

OpenStudy (anonymous):

or if you use the point (-17 , 8) you get y-8 = -1/2 (x-(-17)) which is y-8 = -1/2 (x+17)

OpenStudy (anonymous):

The coordinates of point one are x1,y1 and that of point 2 are x2,y2.So according to formula u use coordinates of first point.Ans Is C then.

OpenStudy (anonymous):

Thank you . . .

OpenStudy (anonymous):

U Are Most Welcome.